OpenStudy (anonymous):

Find an equation for the line with slope m through the point (a,c). ***I'm just forgetting some of the small steps... Could someone please refresh me? I know you start off with y=mx+b... I just forgot how to create the equation when you're using variables! THank you!!!:)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ a}}\quad ,&{\color{blue}{ c}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}} \\ \quad \\ y-y_1={\color{green} m}(x-x_1)\) plug and chug, and solve for "y"

OpenStudy (anonymous):

ohhh so you're using point slope form for this one?

OpenStudy (anonymous):

Well lets start from the easiest part - slope - this is the multiplier of the x As such your equation becomes y = mx + T where t we need to find out. since the line passes through point (a,c) then you know that when x = a, y =c as such c = ma + t - all thats left is solve for T

OpenStudy (anonymous):

so y-c=m(x-a) ?

OpenStudy (jdoe0001):

yes, to get the equation, yes, you'd use the point-slope form

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so do i distribute now?

OpenStudy (jdoe0001):

and you can solve that for "y" to set to slope-intercept form

OpenStudy (anonymous):

okay, so now it's y-c=mx-ma ?

OpenStudy (anonymous):

@Garm when you say t, do you mean the y-int? :/

OpenStudy (anonymous):

T is just unknown like x

OpenStudy (anonymous):

solving for t and then substituting it into previous equation will yield an answer

OpenStudy (anonymous):

ohhh i see :) and @jdoe0001 how do we solve for the "y" ?

OpenStudy (jdoe0001):

add "c" to both sides

OpenStudy (anonymous):

turn formula into something that looks like y=.....

OpenStudy (anonymous):

okay, just want to make sure i'm following correctly so far :) are these the correct steps so far? y-y1=m(x-x1) y-c=m(x-a) y-c=mx-ma +c +c = y=mx-ma+c ??

OpenStudy (anonymous):

did I do that last part right? :/

OpenStudy (jdoe0001):

yes

OpenStudy (anonymous):

okay awesome!! :) do i do something after that? or is that the final equation? :/

OpenStudy (anonymous):

well you know m a and c as such you have nothing else unknown in your equation but x and y - therefore it is a valid formula to draw a line =)

OpenStudy (anonymous):

essentially that is what you are looking for - you need to make sure you only have 2 unknowns for function

OpenStudy (anonymous):

how do i do that? so from y=mx+ma+c, what do I do?

OpenStudy (anonymous):

well what else you dont know in that formula but x and y?

OpenStudy (anonymous):

we don't know m ? a and c?

OpenStudy (anonymous):

you do

OpenStudy (jdoe0001):

did I do that last part right? \(\large \checkmark\)

OpenStudy (anonymous):

m a and c were given to you in the question

OpenStudy (anonymous):

as such they are not unknowns they are essentially just abstract numbers

OpenStudy (anonymous):

ohhh haha okay but y=mx-ma+c has to be simplified more? am I understanding that right? or is that the final equation?

OpenStudy (anonymous):

@jdoe0001 yes all your stuff is top notch

OpenStudy (anonymous):

^^agreed :P and i love the latex! i can never remember how to do the latex stuff haha

OpenStudy (anonymous):

@iheartfood nope! =) that's the beauty of algebra - you dont need to know the answer as number to actually have correct answer.

OpenStudy (anonymous):

awesome!! so the problem is done?! :O it's just y=mx-ma+c ?

OpenStudy (anonymous):

indeed - because you cannot simplify it further.

OpenStudy (anonymous):

okay great!!! :) thank you so much!! @jdoe0001 @Garm I had totally forgotten how to do this with variables and this cleared it up! Thanks again :)

OpenStudy (jdoe0001):

yw